TRANSFERABILITY OF PULAY’S SCALE FACTORS IN THE IVa GROUP OF THE MENDELEYEV PE-RIODIC SYSTEM

\underline{YURII N. PANCHENKO}, Laboratory of Molecular Spectroscopy, Division of Physical Chemistry, Department of Chemistry, M. V. Lomonosov Moscow State University, Moscow 119899, Russian Federation; GEORGE R. DE MARÉ, Laboratoire de Chimie Physique Moléculaire, Faculté des Sciences, CP 160/09, Université Libre de Bruxelles, 50, Av. F. D. Roosevelt, B-1050 Brussels, Belgium; and ALEKSANDER V. ABRAMENKOV, Laboratory of Molecular Structure and Quantum Mechanics, Division of Physical Chemistry, Department of Chemistry, M. V. Lomonosov Moscow State University, Moscow 119899, Russian Federation.

Ab initio quantum mechanical calculations were performed for structures and force fields (HF/6-31G^{∗} // HF/6-31G^{∗} ^{a}) of 3,3-dimethylbutene-1, cyclopropene, 1-methylcyclopropene, and 1-trimethylsilyl-, 1,2-bis(trimethylsilyl)-, 1trimethylgermyl-, 1,2-bis(trimethylgermyl)-, 1-trimethylstannyl-, and 1,2-bis(trimethylstannyl)-3,3-dimethylcyclopropene. Scale factors for correction of the quantum mechanical force fields of cyclopropene, 1-methylcyclopropene, and 3,3dimethylbutene-1 were determined using Pulay’s scaling method.^{b} Only the experimental vibrational frequencies of the light isotopomers of these molecules were used in the scaling procedure. The set of scale factors obtained was transferred to the quantum mechanical force fields of all the other molecules mentioned above. The vibrational problems for these molecules were solved. Complete vibrational analyses were carried out for the whole set of these related compounds. Transferability of scale factors for series of related compounds of cyclopropene with heteroatoms from the IVa group of the Mendeleyev Periodic System of chemical elements was demonstrated.

aP. C. Hariharan and J. A. Pople, Chem. Phys. Lett. \underline{16}, 217 (1972).

bYu. N. Panchenko, P. Pulay and F. Török, J. Mol. Structure \underline{34}, 283 (1976); V. I. Pupyshev, Yu. N. Panchenko, Ch. W. Bock and G. Pongor, J. Chem. Phys. \underline{94}, 1247 (1991); Yu. N. Panchenko, G. R. De Maré and V. I. Pupyshev, J. Phys. Chem. \underline{99}, 17544 (1995); Yu. N. Panchenko, Moscow Univ. Chem. Bull. \underline{51} (5), 23 (1996).